OpenStudy (unklerhaukus):

Solve the following differential equation \[(1+x^2)y''=1\]It is y-absent

OpenStudy (anonymous):

\[(1+x^2)y''=1 \rightarrow y'' = 1/(1+x^2)\] So just integrate twice

OpenStudy (ash2326):

( 1+x^2) y''=1 y''= 1/(+x^2) \[y'=\tan^{-1} x + C1\] C1 is constant of integration integrating above with respect to x \[y= x \tan^{-1} x- 1/2( \ln (x^2+1)) + C1 x+C2\]

OpenStudy (unklerhaukus):

THANKYOUS